# Revisiting English Winogender Schemas for Consistency, Coverage, and Grammatical Case

### v1

@@ Abstract
While measuring bias and robustness in coreference resolution are important goals, such measurements are only as good as the tools we use to measure them with. Winogender schemas (Rudinger et al., 2018) are an influential dataset proposed to evaluate gender bias in coreference resolution, but a closer look at the data reveals issues with the instances that compromise their use for reliable evaluation, including treating different grammatical cases of pronouns in the same way, violations of template constraints, and typographical errors. We identify these issues and fix them, contributing a new dataset: Winogender 2.0. Our changes affect performance with state-of-the-art supervised coreference resolution systems as well as all model sizes of the language model FLAN-T5, with F1 dropping on average 0.1 points. We also propose a new method to evaluate pronominal bias in coreference resolution that goes beyond the binary. With this method and our new dataset which is balanced for grammatical case, we empirically demonstrate that bias characteristics vary not just across pronoun sets, but also across surface forms of those sets.
@@ 1 Introduction
Winogender schemas (Rudinger et al., 2018) are a challenge set proposed to evaluate gender bias in English coreference resolution systems. The dataset has become popular as a result of its careful construction; it has been translated to other languages (Hansson et al., 2021; Stanovsky et al., 2019) and used in framings beyond the original paper’s focus on coreference resolution, e.g., to evaluate natural language inferences (Poliak et al., 2018), intrinsic bias in language models (Kurita et al., 2019), and pronoun fidelity (Gautam et al., 2024).
However, a closer look at the original dataset reveals weaknesses that compromise its use for reliable evaluation, including: treating different grammatical cases of pronouns as equivalent, templates that vary more than just pronouns when they should not, and inconsistencies in the structural constraints they are meant to conform to.
In this paper, we identify issues with the original dataset and fix them to create a new, expanded Winogender 2.0 (§3). [^1] We publicly release our dataset and code at https://github.com/uds-lsv/winogender-2.0. We then empirically show how our fixes affect coreference resolution system performance (§4), and contribute a novel method to evaluate pronominal bias in coreference resolution (§5) that goes beyond the binary and focuses on linguistic gender without mapping it to social gender (Cao and Daumé III, 2021).
Figure 1: Performance on our new Winogender 2.0 dataset with supervised coreference resolution systems (CAW-coref and LingMess), and language models fine-tuned for coreference resolution (SpanBERT) and prompted zero-shot (FLAN-T5).
Winogender 2.0 is more difficult than Winogender 1.0 for state-of-the-art supervised coreference resolution systems as well as large language models untrained for the task. We find that grammatical case, which we balance for in Winogender 2.0, matters for both performance and bias results; accusative pronouns are harder to resolve than nominative or possessive pronouns, and system pronominal bias is not always consistent across different grammatical cases of the same pronoun set.
@@ 2 Background: Winogender Schemas
In the original dataset, the entities consist of 60 occupations paired with 60 matching participants (e.g., accountant is paired with taxpayer; see Appendix B for a full list). A pair of sentence templates is created for each occupation and participant, resulting in 120 templates. Each template can be instantiated with three pronoun sets–he, she, and singular they, for a total of 120 x 3 = 360 sentences.
3 Winogender 2.0 Dataset
Although Winogender schemas are established in the coreference resolution literature, we find issues in the dataset that compromise its use for reliable evaluation. We motivate and describe our fixes for these issues here and Appendix F. Winogender 2.0 uses the same 60 occupation-participant pairs, but contains 240 more templates to cover missing grammatical cases and a neopronoun set xe), for a total of 360 templates x 4 pronouns = 1,440 sentences, which we use for evaluation in §4 and §5.
Grammatical case 1.0 2.0
@@
Table 1: Number of templates per grammatical case in Winogender 1.0 and 2.0 (ours).
@@ 3.1 Issues and Solutions
We hypothesize that systems have different performance and bias characteristics with pronouns in different grammatical cases. [^2] Here, we mean the surface form of the pronoun. However, as Table 1 shows, Winogender 1.0 has a variable number of pronouns per grammatical case, and treats them all as equivalent. To enable more granular evaluation, we balance this distribution in Winogender 2.0.
@@
When evaluating large language models that have not explicitly been trained for coreference resolution on a dataset like Winogender schemas, poor performance could mean that the model simply cannot perform the task (with a given prompt). In its current form, Winogender 1.0 does not allow us to disentangle why bad model performance is bad. For a simple setting to test this (and a useful baseline for all systems), we create single-entity sentences that are parallel to the traditional, more complex double-entity sentences.
Version Data LingMess CAW-coref SpanBERT FLAN-T5
@@
Table 2: F1 scores of coreference resolution systems on Winogender 1.0 and Winogender 2.0. In addition to total F1, we report performance split by grammatical case, and F1 consistency on templates grouped by pronouns (PronounC) and grouped by disambiguating pairs (DisambigC). Red, italicized numbers are worse than chance.
@@
For a controlled evaluation comparing pronouns, it is common to use templates that only vary the pronoun. However, 17% of Winogender 1.0 templates must be modified to work with singular they due to its different verbal agreement (“he was” but “they were”). To ensure a fair comparison between pronouns, we modify these templates to work with any pronouns, including neopronouns (Lauscher et al., 2022).
@@
Winograd-like schemas have strict structural constraints so that models cannot inflate performance through heuristics. However, when analyzing Winogender 1.0, we found violations of these constraints, e.g., non-parallel paired templates. We fixed these along with typographical errors to ensure robust and reliable evaluation.
@@ 4 Performance and Consistency
To demonstrate the effects of our changes, we evaluate Winogender 1.0 and 2.0 with a range of models, including supervised coreference resolution systems and large language models.
@@ 4.2 Performance results
Table 2 shows how model performance changes with our new dataset. Firstly, we find that Winogender 2.0 is consistently harder than Winogender 1.0 across all systems, with performance dropping on average by 0.1 F1 points. Scaling behaviour generally holds for SpanBERT and FLAN-T5. In fact, scale seems to supercede supervision, as the largest FLAN-T5 models perform the best overall.
@@ 4.3 Consistency results
We follow Ravichander et al. (2022) in evaluating system robustness via the consistency of outputs across groups of closely related instances, i.e., the score of the lowest-performing instance in the group is taken as the group’s score. We consider two groups: (a) pronoun consistency, and (b) disambiguation consistency. Pronoun consistency measures whether a template can be solved regardless of the pronoun it is filled with. Since we consider four pronoun sets, chance is $0.5^{4}$, or $0.0625$. Inspired by pair accuracy on Winograd schemas  (Abdou et al., 2020), we also measure disambiguation consistency on Winogender 2.0. Here, we keep the pronoun fixed and consider the template pairs that disambiguate the antecedent as either the occupation or the participant (sentences (a) and (b) in Figure 2). Chance is therefore $0.5^{2}$, or $0.25$.
@@ 6 Conclusion
We demonstrate a number of issues with the well-known Winogender schemas dataset, which we fix in our new, expanded Winogender 2.0 dataset. In addition, we propose a novel way to evaluate pronominal bias in coreference resolution that goes beyond the binary and focuses on lexical gender. With our new dataset and method, we evaluate both supervised coreference resolution systems and language models, and find that our changes impact performance and bias characteristics. Our work demonstrates that measurements of bias and robustness are only as good as the datasets and metrics we use to measure them, and we call for careful attention when developing resources for and discussing performance on coreference resolution.
@@
As in Winogender 1.0, our schemas are not “Google-proof” and could conceivably be solved with heuristics, including word co-occurrences, which is a primary concern when creating and evaluating Winograd schemas (Levesque et al., 2012; Amsili and Seminck, 2017; Elazar et al., 2021). The fact that we do not control for this means that our dataset gives generous estimates of model performance and bias, but it also means that this dataset is inappropriate to test model “reasoning.” Our dataset construction instead controls for simple system heuristics that are relevant for coreference resolution, such as always picking the first entity in the sentence, or always picking the second. In interpreting system performance, we also echo Rudinger et al.’s (2018) view of Winogender schemas as having “high positive predictive value and low negative predictive value.” In other words, that they may demonstrate evidence of pronominal bias in systems, but not prove its absence.
We take steps to prevent data contamination following Jacovi et al. (2023), including not releasing our data in plain text, and not evaluating with models behind closed APIs that do not guarantee that our data will not be used to train future models (Balloccu et al., 2024). However, as we cannot guarantee a complete absence of data leakage unless we never release the dataset, we encourage caution in interpreting results on Winogender 2.0 with models trained on data after August 2024.
@@ Acknowledgements
The authors thank Timm Dill for several rounds of patient annotation. Vagrant Gautam received funding from the BMBF’s (German Federal Ministry of Education and Research) SLIK project under the grant 01IS22015C. Anne Lauscher’s work is funded under the Excellence Strategy of the German Federal Government and States.
@@ Appendix G Additional Results
Table 5: F1 scores of coreference resolution systems on Winogender 2.0, split by pronoun series. Red, italicized numbers are worse than chance.
@@
Table 6: F1 scores of coreference resolution systems on single-entity sentences in Winogender 2.0. In addition to total F1, we report performance split by grammatical case and by pronoun series.

### v2

# WinoPron: Revisiting English Winogender Schemas for Consistency, Coverage, and Grammatical Case

@@ Abstract
While measuring bias and robustness in coreference resolution are important goals, such measurements are only as good as the tools we use to measure them with. Winogender schemas (Rudinger et al., 2018) are an influential dataset proposed to evaluate gender bias in coreference resolution, but a closer look reveals issues with the data that compromise its use for reliable evaluation, including treating different pronominal forms as equivalent, violations of template constraints, and typographical errors. We identify these issues and fix them, contributing a new dataset: WinoPron. Our changes affect performance with state-of-the-art supervised coreference resolution systems as well as all model sizes of the language model FLAN-T5, with F1 dropping on average 10 percentage points. We also propose a new method to evaluate pronominal bias in coreference resolution that goes beyond the binary. With this method and our new dataset which is balanced for grammatical case, we empirically demonstrate that bias characteristics vary not just across pronoun sets, but also across surface forms of those sets.
@@ 1 Introduction
Third-person pronouns (he, she, they, etc.) help us refer to people in conversation. Since they mark referential gender in English, gender bias affects how coreference resolution systems map these pronouns to people. Rudinger et al. (2018) demonstrated this by introducing Winogender schemas, a challenge dataset to evaluate occupational gender bias in coreference resolution systems. The dataset has become popular due to its careful construction; it has been translated to other languages (Hansson et al., 2021; Stanovsky et al., 2019) and used in framings beyond coreference resolution, e.g., to evaluate natural language inferences (Poliak et al., 2018), intrinsic bias in language models (Kurita et al., 2019), and pronoun fidelity (Gautam et al., 2024).
However, a closer look at the dataset reveals weaknesses that compromise its use for reliable evaluation, including treating different pronominal forms as equivalent, varying more than just pronouns in templates that should otherwise be identical, and not conforming to structural requirements.
In this paper, we identify issues with the original dataset and fix them to create a new, expanded WinoPron3). [^1] Data and code available at github.com/uds-lsv/winopron. We then empirically show how our fixes affect coreference resolution system performance (§4), and contribute a novel method to evaluate pronominal bias in coreference resolution (§5) that goes beyond the binary and focuses on linguistic gender without mapping it to social gender (Cao and Daumé III, 2021).
Figure 1: Performance on our new WinoPron dataset with supervised coreference resolution systems (CAW-coref and LingMess), and language models fine-tuned for coreference resolution (SpanBERT) and prompted zero-shot (FLAN-T5).
WinoPron is more difficult than Winogender schemas for state-of-the-art supervised coreference resolution systems as well as large language models untrained for the task. We find that grammatical case, which we balance for in WinoPron, matters for both performance and bias results; accusative pronouns are harder to resolve than nominative or possessive pronouns, and system pronominal bias is not always consistent across different grammatical cases of the same pronoun set.
@@ 2 Background: Winogender Schemas
The entities consist of 60 occupation-participant pairs (e.g., accountant is paired with taxpayer; see Appendix B for a full list). A pair of templates is created for each occupation and participant, resulting in 120 templates. Each template can be instantiated with three pronoun sets (he, she, and singular they), for a total of 120 x 3 = 360 sentences.
3 WinoPron Dataset
Although Winogender schemas are established in the coreference resolution literature, we find issues in the dataset that compromise its use for reliable evaluation. We motivate them and describe our fixes here and in Appendix F, creating WinoPron. We reuse the occupation-participant pairs from Winogender schemas, but add 240 templates to cover missing grammatical cases for a total of 360 templates. We also include a neopronoun set (xe), giving us 360 templates x 4 pronouns = 1,440 sentences, which we use for evaluation in §4 and §5.
Grammatical case WS WP
@@ 3 WinoPron Dataset
Table 1: Number of templates per grammatical case in Winogender schemas (WS) and WinoPron (WP).
@@ 3.1 Issues and Solutions
We hypothesize that systems have different performance and bias characteristics with pronouns in different grammatical cases. [^2] Here, we mean the surface form of the pronoun. However, as Table 1 shows, Winogender schemas have a variable number of pronouns per grammatical case, and treat them all as equivalent. To enable more granular evaluation, we balance this distribution in WinoPron.
@@
When evaluating large language models on coreference resolution when they have not explicitly been trained for it, poor performance could mean that the model simply cannot perform the task (with a given prompt). In its current form, Winogender schemas do not allow us to disentangle why bad model performance is bad. In WinoPron, we create single-entity sentences that are parallel to the traditional, more complex double-entity sentences, for a simple setting to test this, and a useful baseline for all systems.
Dataset Data LingMess CAW-coref SpanBERT FLAN-T5
@@
Table 2: F1 scores of coreference resolution systems on Winogender schemas (WS) and WinoPron (WP). We report total F1, performance split by grammatical case, and F1 consistency on templates grouped by pronouns (PronounC) and grouped by disambiguating pairs (DisambigC). Red, italicized numbers are worse than chance.
@@
For a controlled evaluation comparing pronouns, it is common to use templates that only vary the pronoun. However, 17% of Winogender schemas must be modified to work with singular they due to its different verbal agreement (“he was” but “they were”). To ensure a fair comparison between pronouns, we modify these templates to work with any pronouns, including neopronouns (Lauscher et al., 2022).
@@
Winograd-like schemas have strict structural constraints so that models cannot inflate performance through heuristics. However, when analyzing Winogender schemas, we found constraint violations, e.g., non-parallel paired templates. We fixed these along with typographical errors to ensure robust and reliable evaluation.
@@ 4 Performance and Consistency
To demonstrate the effects of our changes, we evaluate Winogender schemas and WinoPron with a range of models, including supervised coreference resolution systems and large language models.
@@ 4.2 Performance results
Table 2 shows how model performance changes with our new dataset. Firstly, WinoPron is consistently harder than Winogender schemas across all systems, with F1 dropping on average by 10 percentage points. Scaling behaviour generally holds for SpanBERT and FLAN-T5. In fact, scale seems to supercede supervision, as the largest FLAN-T5 models perform the best overall.
@@ 4.3 Consistency results
We follow Ravichander et al. (2022) in evaluating system robustness via the consistency of outputs across groups of closely related instances, i.e., the score of the lowest-performing instance in the group is taken as the group’s score. We consider two groups: (a) pronoun consistency, and (b) disambiguation consistency. Pronoun consistency measures whether a template can be solved regardless of the pronoun it is filled with. Since we consider four pronoun sets, chance is $0.5^{4}$, or $0.0625$. Inspired by pair accuracy on Winograd schemas  (Abdou et al., 2020), we also measure disambiguation consistency on WinoPron. Here, we keep the pronoun fixed and consider the template pairs that disambiguate the antecedent as either the occupation or the participant (sentences (a) and (b) in Figure 2). Chance is therefore $0.5^{2}$, or $0.25$.
@@ 6 Conclusion
We demonstrate a number of issues with the well-known Winogender schemas dataset, which we fix in our new, expanded WinoPron dataset. In addition, we propose a novel way to evaluate pronominal bias in coreference resolution that goes beyond the binary and focuses on lexical gender. With our new dataset and method, we evaluate both supervised coreference resolution systems and language models, and find that our changes impact performance and bias characteristics. Our work demonstrates that measurements of bias and robustness are only as good as the datasets and metrics we use to measure them, and we call for careful attention when developing resources for and discussing performance on coreference resolution.
@@
As in Winogender schemas, our schemas are not “Google-proof” and could conceivably be solved with heuristics, including word co-occurrences, which is a primary concern when creating and evaluating Winograd schemas (Levesque et al., 2012; Amsili and Seminck, 2017; Elazar et al., 2021). The fact that we do not control for this means that our dataset gives generous estimates of model performance and bias, but it also means that this dataset is inappropriate to test model “reasoning.” Our dataset construction instead controls for simple system heuristics that are relevant for coreference resolution, such as always picking the first entity in the sentence, or always picking the second. In interpreting system performance, we also echo Rudinger et al.’s (2018) view of Winogender schemas as having “high positive predictive value and low negative predictive value.” In other words, that they may demonstrate evidence of pronominal bias in systems, but not prove its absence.
We take steps to prevent data contamination following Jacovi et al. (2023), including not releasing our data in plain text, and not evaluating with models behind closed APIs that do not guarantee that our data will not be used to train future models (Balloccu et al., 2024). However, as we cannot guarantee a complete absence of data leakage unless we never release the dataset, we encourage caution in interpreting results on WinoPron with models trained on data after August 2024.
@@ Acknowledgements
The authors thank Timm Dill for several rounds of patient annotation, and are grateful to Rachel Rudinger and Benjamin Van Durme for their comments. Vagrant Gautam received funding from the BMBF’s (German Federal Ministry of Education and Research) SLIK project under the grant 01IS22015C. Anne Lauscher’s work is funded under the Excellence Strategy of the German Federal Government and States.
@@ Appendix G Additional Results
Table 5: F1 scores of coreference resolution systems on WinoPron, split by pronoun series. Red, italicized numbers are worse than chance.
@@
Table 6: F1 scores of coreference resolution systems on single-entity sentences in WinoPron. In addition to total F1, we report performance split by grammatical case and by pronoun series.